OpenStudy (anonymous):

What is the sum of __3 2x and __7 4x

OpenStudy (smurfy14):

can u rewrite it i dont really get how u wrote it

OpenStudy (anonymous):

the sum of 3over 2x and 7 over 4x \[3\]

OpenStudy (saifoo.khan):

13/4x

OpenStudy (smurfy14):

13 over 4x

OpenStudy (anonymous):

3/2x + 7/4x we will take LCM of denominator (6+7)/4x 13/4x Answer is 13/4x

OpenStudy (smurfy14):

\[13 \over 4x\]

OpenStudy (anonymous):

(3/2x) * (2/2) = 6/4x 6/4x + 7/4x = 13/4x

OpenStudy (anonymous):

Yes. 13 over 4x